OpenStudy (anonymous):

find the missing value to the nearest hundredth. sin _______ = 8/15

OpenStudy (anonymous):

57.77 32.23 28.07 27.82

OpenStudy (jdoe0001):

\(\bf sin(\theta)=\cfrac{8}{15}\implies sin^{-1}[sin(\theta)]=sin^{-1}\left(\cfrac{8}{15}\right)\implies \theta=sin^{-1}\left(\cfrac{8}{15}\right)\)

OpenStudy (anonymous):

huh